A stock's returns are normally distributed with a mean of 8% pa and a standard deviation of 15 percentage points pa. What is the 99% confidence interval of returns over the next year? Note that the Z-statistic corresponding to a one-tail: • 90% normal probability density function is 1.282. • 95% normal probability density function is 1.645. • 97.5% normal probability density function is 1.960. • 99% normal probability density function is 2.326. • 99.5% normal probability density function is 2.576.
Answer #1

µ = 8

sd = 15

Z for 99% confidence interval = Z0.005 = 2.576

Confidence interval = (µ + Z0.005 * sd)

                                = (8 + 2.576 * 15)

                                = (8 + 38.64)

                                = (-30.64% , 46.64%)